OpenStudy (anonymous):

Can someone please help me? WILL FAN AND GIVE MEDAL Linda is studying the sale of a particular brand of cereals from the year 1993 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 5t + 52 Part A: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work.

OpenStudy (error1603):

k

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A. The sale in 1993 B. domain = [1993, 2004] C. first year = 1993, sale = S(0) = 52 4th year = 1997, sale = S(4) = 88 the average rate of change from 1993 to 1997 is r = (S(4) - S(0))/4 = (88 - 52)/4 = 36/4 = 9 million $ per year if the first year is 1994 and the fourth 1998 then r = (S(5) - S(0))/4 = 11 million $ per year

OpenStudy (anonymous):

Did you copy and paste that from Answer.com??